# Energetic Resilience of Linear Driftless Systems

### v2

@@ III Nominal and Malfunctioning Energies
We wish to find the solution of the following problem:
$\operatorname*{arg\,min}_{z}$ $\int_{0}^{t_{f}}z^{T}(t)z(t)\mathrm{d}t$
[…]
for a given $\overline{z}\in\mathbb{R}^{n_{z}}$. This is a constrained calculus of variations problem [23]. Let $L(t,z)=z^{T}(t)z(t)$ and $M(t,z)=z(t)$. The first-order necessary condition for optimality states that the solution $z^{*}(t)$ must satisfy the Euler-Lagrange equation for $\lambda_{0}^{*}L(t,z)+\lambda^{*T}M(t,z)$ for $(\lambda_{0}^{*},\lambda^{*})\neq(0,0)$ [23]. Thus,
${\frac{\partial}{\partial z}\left(\lambda_{0}^{*}z^{*T}(t)z^{*}(t)+\lambda^{*T}z^{*}(t)\right)}$
$=$ ${\frac{d}{dt}\left[\frac{\partial}{\partial z^{\prime}}\left(\lambda_{0}^{*}z^{*T}(t)z^{*}(t)+\lambda^{*T}z^{*}(t)\right)\right]=0,}$
where the last equality holds since $L(t,z)$ and $M(t,z)$ are independent of $z^{\prime}=\frac{dz}{dt}$. Then, $z^{*}(t)$ satisfies
$\lambda_{0}^{*}z^{*}(t)+\lambda^{*}=0.$
If $\lambda_{0}^{*}=0$, then $\lambda^{*}=0$, contradicting the requirement that $(\lambda_{0}^{*},\lambda^{*})\neq(0,0)$. Thus, $\lambda_{0}^{*}\neq 0$, and
$z^{*}(t)=-\frac{\lambda^{*}}{\lambda_{0}^{*}},$
i.e. a constant. Given the constraint $\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)=\overline{z}$, we have $z^{*}(t)=\overline{z}\leavevmode\nobreak\ \text{ for all $t\in[0,t_{f}]$.}$
@@
III-A Nominal Energy
Consider the system (1). Solving for $x(t_{f})=0$ from these dynamics,
@@ III-A Nominal Energy
Since we wish to find the minimum energy control signal $u$, we require the solution $\overline{u}$ with the least norm $\|\overline{u}\|_{2}$, based on Lemma 1. This least-norm solution [24, Corollary 2] is given by
@@
where $B^{\dagger}$ is defined as in [20]. As the nominal dynamics (1) are assumed to be controllable, a control law achieves finite-time stabilization if and only if its mean value over the interval $[0,t_{f}]$ is given by (13). Now, we recall the constraint $u\in\operatorname*{\mathcal{U}}$, i.e., $\|u(t)\|_{\infty}\leq 1$ for all $t$. For any control law $u(t)$ with mean value $\overline{u}^{LS}$, the condition $\|\overline{u}^{LS}\|_{\infty}\leq 1$ is necessary to ensure $\|u(t)\|_{\infty}\leq 1$ for all $t$. If this were not true, there would exist at least one component $i$ of the control, denoted $u_{i}(t)$, and at least one instant of time $t_{1}$ where $u_{i}(t_{1})\geq 1$, violating the constraint $u\in\operatorname*{\mathcal{U}}$. Next, $\|\overline{u}^{LS}\|_{\infty}\leq 1$ if and only if
@@
Thus, condition (14) is necessary to ensure that a control law achieving finite-time stabilization also satisfies the constraint $u\in\operatorname*{\mathcal{U}}$.
Note from Lemma 1 that over all signals with a given mean value, the minimum energy signal is a constant equal to that mean value. Thus, the signal achieving the infimum in (5), denoted $u^{*}(t)$, is
$u^{*}(t)=\overline{u}^{LS}=-\frac{1}{t_{f}}B^{\dagger}x_{0}\leavevmode\nobreak\ \text{ for all }\leavevmode\nobreak\ t\in[0,t_{f}].$ (15)
The nominal energy is the energy of this optimal control signal, and thus
@@
Since $u^{*}(t)=\overline{u}^{LS}$, condition (14) is both necessary and sufficient to ensure $u^{*}\in\operatorname*{\mathcal{U}}$. Equations (15) and (16), along with condition (14), are the central results of this subsection.
III-B Malfunctioning Energy
We now consider the malfunctioning system (3) and derive expressions for the malfunctioning (6) and worst-case total (8) energies. Writing out the solution to (3) and setting $x(t_{f})=0$, we have
@@ III-B Malfunctioning Energy
Checking this condition is equivalent to checking
$\max_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\frac{1}{t_{f}}\left[B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right]_{i}\leq 1$
and $\max_{\|\overline{u}_{uc}\|_{\infty}\leq 1}-\frac{1}{t_{f}}\left[B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right]_{i}\leq 1,$
for every $i$, where the subscript $i$ denotes the $i$-th component of the vector $B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\in\mathbb{R}^{m}$. The above conditions are all linear in $\overline{u}_{uc}$, and thus it is sufficient to check these conditions on the vertices of the hypercube $\|\overline{u}_{uc}\|_{\infty}\leq 1$, which can be done efficiently [25]. Finally, as a consequence of Lemma 1, we note that the control signal $u_{c}^{*}$ achieving the infimum in (6) is a constant equal to the mean value $\overline{u}_{c}^{LS}$. Thus, under the condition (19) on $t_{f}$,
$u_{c}^{*}(t)=\overline{u}_{c}^{LS}=-\frac{1}{t_{f}}B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ t\in[0,t_{f}].$ (20)
The corresponding malfunctioning energy is
@@
Note that the optimal controlled input $u_{c}^{*}$ in (20) and the malfunctioning energy $E_{M}^{*}(x_{0},t_{f},u_{uc})$ in (21) are both dependent on the uncontrolled input $u_{uc}$. To quantify the maximal effect of $u_{uc}$, we now consider the worst-case total energy defined in (8). Using (21), we have
@@
$\hskip 39.83368pt+\int_{0}^{t_{f}}u_{uc}^{T}(t)u_{uc}(t)\mathrm{d}t\Biggr{\}}.$ (22)
A closed-form, analytical expression for the above problem is not straightforward to obtain, and hence we focus on bounding these terms. First, note that
$\overline{E}_{M}(x_{0},t_{f})$ $\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}t_{f}\overline{u}_{uc}^{T}B_{uc}^{T}B_{uc}\overline{u}_{uc}}_{=T_{1}}$
$+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}2x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}}_{=T_{2}}$
$+\underbrace{\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\int_{0}^{t_{f}}u_{uc}^{T}(t)u_{uc}(t)\mathrm{d}t}_{=T_{3}},$ (23)
where the inequality is obtained on splitting the supremum over three different terms. We first consider $T_{1}$. Note that $B_{uc}^{T}B_{uc}\in\mathbb{R}^{p\times p}$ is positive semi-definite. Let $B_{uc}^{T}B_{uc}=V\Lambda V^{T}$ be its spectral decomposition where $\Lambda=\operatorname*{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$ with each $\lambda_{i}\geq 0$ and $V$ is the orthonormal matrix of eigenvectors of $B_{uc}^{T}B_{uc}$. Further, let $q=V^{T}\overline{u}_{uc}$. Then, $\|q\|_{\infty}\leq\|V^{T}\|_{\infty}\|\overline{u}_{uc}\|_{\infty}\leq\|V\|_{1}$, where we use the sub-multiplicative property of norms, $\|V^{T}\|_{\infty}=\|V\|_{1}$ and $\|\overline{u}_{uc}\|_{\infty}\leq 1$. Then,
@@
where the optimal $q^{*}$ has each component $q_{i}^{*}=\pm\|V\|_{1}$, maximizing the objective since each $\lambda_{i}>0$. The upper bound in (24) is obtained by replacing the set $\|\overline{u}_{uc}\|_{\infty}\leq 1$ by the set $\|q\|_{\infty}=\|V^{T}\overline{u}_{uc}\|\leq\|V\|_{1}$. Due to the use of the sub-multiplicative property, the latter set is larger. Next, consider $T_{2}$. For any vector $\eta\in\mathbb{R}^{p}$, $\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\eta^{T}\overline{u}_{uc}=\|\eta\|_{1},$ where the optimal $\overline{u}_{uc}$ is given by $\overline{u}_{uc}^{*}=\operatorname*{sign}(\eta)$. Thus,
@@
Finally, using (23), (24), (25) and (26), we have
$+2\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\|_{1}.$ (27)
Remark 1.
@@
The bound obtained in (27) may be quite conservative. However, in Section IV, we consider the special case when $p=1$, i.e., when control authority is lost over a single actuator, and show that a closed-form expression for $\overline{E}_{M}(x_{0},t_{f})$ can be derived. We then use that expression to derive bounds on the energetic resilience metrics (9) and (10). These metrics compare the nominal energy $E_{N}^{*}(x_{0},t_{f})$ and the worst-case total energy $\overline{E}_{M}(x_{0},t_{f})$.
Remark 2.
@@
We note some similarities between the results of this section and the works of Bouvier et al. [14, 26], which discuss time-optimality for linear driftless systems. We have shown that the energy-optimal control signals in the nominal (15) and malfunctioning (20) cases are constant in the interval $[0,t_{f}]$. Similarly, it was shown in [14] that the time-optimal control signals in both the nominal and malfunctioning cases are constant. We also note in the malfunctioning case that each component of the worst-case uncontrolled input $u_{uc}^{*}(t)$ takes on a constant value of either $+1$ or $-1$ in the interval $[0,t_{f}]$. A similar result in [14] shows that the worst-case uncontrolled input to maximize the reachable time also has constant components with the maximum allowable amplitude, using an optimization result derived in [26]. A contrasting feature of our work is that we consider energy-optimality as our objective, while the work in [14] considers time-optimality. Further, we are able to derive closed-form expressions for the optimal control in both the nominal (15) and malfunctioning (20) cases, which is not accomplished in [14]. We also derive a closed-form expression for the worst-case uncontrolled input $u_{uc}^{*}(t)$, further discussed in Section IV. Such an expression is not provided in [14].
III-C Final Time Design
We now briefly consider a problem of final time design for the malfunctioning system. Given that the optimal control signal is used in the malfunctioning case, we wish to find the optimal final time $t_{f}^{*}$ that minimizes the malfunctioning energy for a given uncontrolled input $u_{uc}$. This is useful in designing control signals for tasks with a variable completion time and constraints on the total available control energy. In other words, we wish to solve the following problem:
$t_{f}^{*}$ $=\operatorname*{arg\,min}_{t_{f}}E_{M}^{*}(x_{0},t_{f},u_{uc})$
$=\operatorname*{arg\,min}_{t_{f}}\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}.$ (28)
Using the first-order stationary condition $\frac{\partial}{\partial t_{f}}\left(E_{M}^{*}(x_{0},t_{f},u_{uc})\right)=0$, we have
$t_{f}\frac{\partial}{\partial t_{f}}\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}=\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2},$
with $t_{f}\neq 0$. The above equality simplifies to the following quadratic equation in $t_{f}$:
$t_{f}^{2}\overline{u}_{uc}^{T}B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}=x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}.$
Solving for $t_{f}>0$, we have:
$t_{f}^{*}=\frac{\|B_{c}^{\dagger}x_{0}\|_{2}}{\|B_{c}^{\dagger}B_{uc}\overline{u}_{uc}\|_{2}}.$ (29)
It is then straightforward to show that
$\frac{\partial^{2}}{\partial t_{f}^{2}}\left(E_{M}^{*}(x_{0},t_{f},u_{uc})\right)=\frac{2\|B_{c}^{\dagger}x_{0}\|_{2}^{2}}{t_{f}^{3}}\leavevmode\nobreak\ \text{ and thus}\leavevmode\nobreak$
$\frac{\partial^{2}}{\partial t_{f}^{2}}\left(E_{M}^{*}(x_{0},t_{f},u_{uc})\right)\Bigg{|}_{t_{f}=t_{f}^{*}}=\frac{2\|B_{c}^{\dagger}B_{uc}\overline{u}_{uc}\|_{2}^{3}}{\|B_{c}^{\dagger}x_{0}\|_{2}}>0.$
Hence $t_{f}^{*}$ in (29) is the unique minimum for the problem in (III-C). The result (29) thus provides the final time that minimizes the malfunctioning energy for a given uncontrolled input $\overline{u}_{uc}$.
IV Resilience to the Loss of One Actuator
@@
In this section, we consider the special case of losing control authority over one actuator, i.e., $p=1$. We derive a closed-form expression for the worst-case total energy (8) and for the corresponding worst-case uncontrolled input. Using this expression, we derive bounds on the energetic resilience metrics (9) and (10), which compare the worst-case total energy to the nominal energy.
IV-A Worst-case Total Energy
When $p=1$, $B_{uc}\in\mathbb{R}^{n}$ and $u_{uc}(t)\in\mathbb{R}$. Using (21), we can first obtain the total energy in (7) as
$E_{M}^{+}(x_{0},t_{f},u_{uc})$
$\hskip 10.00002pt=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}+\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t.$ (30)
Then, from (8),
@@ IV-A Worst-case Total Energy
$=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}$
$+\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\Biggl{\{}t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}+2x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}$
$+\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t\Biggr{\}}.$ (31)
Recall that $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ implies $\|\overline{u}_{uc}\|_{\infty}\leq 1$, or $-1\leq\overline{u}_{uc}\leq 1$. Considering just the second term inside the supremum, it is easily seen that
$\sup_{-1\leq\overline{u}_{uc}\leq 1}2x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}=2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|,$
where an optimal $\overline{u}_{uc}$ for this particular term is $\overline{u}_{uc}^{*}=\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})$.
If $\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})\in\{-1,+1\}$, i.e., $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\neq 0$, this choice of $\overline{u}_{uc}^{*}$ also maximizes the first term in the supremum in (31), with $t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}=t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{*2}=t_{f}\|B_{uc}\|_{2}^{2}$. Further, the corresponding worst-case uncontrolled input $u_{uc}^{*}(t)$ takes on a constant value of either $+1$ or $-1$ in the interval $[0,t_{f}]$. Such a $u_{uc}^{*}(t)$ maximizes the third term in the supremum, since it achieves the maximum possible value of the integrand $u_{uc}^{2}(t)=1$ at every instant of time, for $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$. The corresponding value of the third term is $t_{f}$. Thus, when $\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})\in\{-1,+1\}$, the constant uncontrolled input
$u_{uc}^{*}(t)=\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})\leavevmode\nobreak\ \text{ for all }\leavevmode\nobreak\ t\in[0,t_{f}]$ (32)
maximizes each term in the supremum in (31), and hence,
@@
If $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}=0$, the second term inside the supremum in (31) vanishes, irrespective of what value $\overline{u}_{uc}$ takes. The first and third terms are still maximized by uncontrolled inputs $u_{uc}^{*}(t)$ that take on a constant value of $+1$ or $-1$ in the interval $[0,t_{f}]$, and the corresponding worst-case total energy is simply given by
$\overline{E}_{M}(x_{0},t_{f})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right).$ (34)
In what follows, we assume $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\neq 0$. Using (33), we derive bounds on the energetic resilience metrics (9) and (10) to quantify how much larger the worst-case total energy $\overline{E}_{M}(x_{0},t_{f})$ is when compared to the nominal energy $E_{N}^{*}(x_{0},t_{f})$.
IV-B Energetic Resilience Metrics
We now consider the energetic resilience metrics $r_{A}(t_{f},R)$ in (9) and $r_{M}(t_{f},R)$ in (10), and use the expression (33) to derive bounds on the these metrics in the case when control authority is lost over a single actuator and when $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\neq 0$. We note that bounds on $r_{A}(t_{f},R)$ and $r_{M}(t_{f},R)$ can be derived when control authority is lost over $p>1$ actuators using (27). However, these bounds are more conservative since (27) is only an upper bound on the worst-case total energy and not an exact expression as in (33). Further, if $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}=0$, the expression (34) for the worst-case total energy can be used to derive bounds on energetic resilience metrics.
Proposition 1.
For the case when $p=1$, the additive energetic resilience metric $r_{A}(t_{f},R)$ is bounded from above as follows:
$r_{A}(t_{f},R)$ $\leq\frac{R^{2}}{t_{f}}\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)$
$+2R\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right).$ (35)
Proof.
Using (16), (33) and (9),
$r_{A}(t_{f},R)$ $=\sup_{\|x_{0}\|_{2}\leq R}\Biggl{\{}\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)$
$+2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|-\frac{1}{t_{f}}\left\|B^{\dagger}x_{0}\right\|_{2}^{2}\Biggr{\}}$
$\leq t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)+2R\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}$
$+\sup_{\|x_{0}\|_{2}\leq R}\frac{1}{t_{f}}x_{0}^{T}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)x_{0},$ (36)
where the inequality follows from the Cauchy-Schwarz inequality applied to the term $2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|$. Note that $B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}$ is symmetric, and hence using the Rayleigh inequality and (36),
$r_{A}(t_{f},R)$ $\leq\frac{R^{2}}{t_{f}}\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)$
$+2R\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right).$
concluding the proof. ∎
Proposition 2.
For the case when $p=1$, the multiplicative metric $r_{M}(t_{f},R)$ is bounded from below as follows:

### v3

@@ 3 Nominal and Malfunctioning Energies
The result follows from the first-order condition for optimality using the Euler-Lagrange equation (Liberzon, 2012), for the constrained calculus of variations problem
$\operatorname*{arg\,min}_{z}\int_{0}^{t_{f}}z^{T}(t)z(t)\mathrm{d}t~{}\text{ s.t. }~{}\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)\mathrm{d}t=\overline{z}.$
@@
3.1 Nominal Energy
Consider the system (1). Solving for $x(t_{f})=0$,
@@ 3.1 Nominal Energy
Since we wish to find the minimum energy control signal $u$, we require the solution $\overline{u}$ with the least norm $\|\overline{u}\|_{2}$, based on Lemma 1. This least-norm solution (Penrose, 1956, Corollary 2) is given by
@@
where $B^{\dagger}$ is defined as in Penrose (1955). As the nominal dynamics (1) are assumed to be controllable, a control law achieves finite-time stabilization if and only if its mean value over the interval $[0,t_{f}]$ is given by (12). Next, note that the condition $\|\overline{u}^{LS}\|_{\infty}\leq 1$ is necessary to ensure $u\in\operatorname*{\mathcal{U}}$. This condition is satisfied if and only if
@@
Thus, condition (13) is necessary to ensure that a control $u$ achieving $x(t_{f})=0$ also satisfies $u\in\operatorname*{\mathcal{U}}$. From Lemma 1, over all signals with a given mean value, the minimum energy signal is a constant equal to that mean value. Thus, the signal achieving the infimum in (5), denoted $u^{*}(t)$, is
$u^{*}(t)=\overline{u}^{LS}=-\frac{1}{t_{f}}B^{\dagger}x_{0}~{}\text{ for all }~{}t\in[0,t_{f}],$ (14)
from which
@@
Since $u^{*}(t)=\overline{u}^{LS}$, condition (13) is both necessary and sufficient to ensure $u^{*}\in\operatorname*{\mathcal{U}}$.
3.2 Malfunctioning Energy
We now consider the malfunctioning system (3) and derive expressions for the malfunctioning (6) and worst-case total (8) energies. Setting $x(t_{f})=0$ in the solution to (3),
@@ 3.2 Malfunctioning Energy
This condition can easily be decomposed into a set of conditions that are linear in $\overline{u}_{uc}$, which can be checked efficiently on the vertices of the hypercube $\|\overline{u}_{uc}\|_{\infty}\leq 1$ (Bertsimas and Tsitsiklis, 1997). Finally, as a consequence of Lemma 1, we note that the control signal $u_{c}^{*}$ achieving the infimum in (6) is a constant equal to the mean value $\overline{u}_{c}^{LS}$. Thus, under the condition (18) on $t_{f}$,
$u_{c}^{*}(t)=\overline{u}_{c}^{LS}=-\frac{1}{t_{f}}B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)~{}\text{for all}~{}t\in[0,t_{f}].$ (19)
Clearly, the corresponding malfunctioning energy is
@@
We now aim to quantify the maximal effect of the uncontrolled input $u_{uc}$ on both the optimal controlled input $u_{c}^{*}$ in (19) and the malfunctioning energy $E_{M}^{*}(x_{0},t_{f},u_{uc})$ in (20), using the worst-case total energy $\overline{E}_{M}(x_{0},t_{f})$ defined in (8). A closed-form, analytical expression for the supremum in (8) using (20) is not straightforward to obtain, and hence we focus on bounding $\overline{E}_{M}(x_{0},t_{f})$.
Proposition 1 (Worst-case Total Energy)
The worst-case total energy $\overline{E}_{M}(x_{0},t_{f})$ is bounded from above as follows:
$+2\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\|_{1}.$ (21)
Proof.
Using (20), we have
@@
$+\int_{0}^{t_{f}}u_{uc}^{T}(t)u_{uc}(t)\mathrm{d}t\Biggr{\}}$
$\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}t_{f}\overline{u}_{uc}^{T}B_{uc}^{T}B_{uc}\overline{u}_{uc}}_{=T_{1}}$
$+\underbrace{\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}2x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}}_{=T_{2}}+\underbrace{\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\int_{0}^{t_{f}}u_{uc}^{T}(t)u_{uc}(t)\mathrm{d}t}_{=T_{3}},$ (22)
where the inequality is obtained on splitting the supremum over three different terms. We first consider $T_{1}$. Note that $B_{uc}^{T}B_{uc}\in\mathbb{R}^{p\times p}$ is positive semi-definite. Let $B_{uc}^{T}B_{uc}=V\Lambda V^{T}$ be its spectral decomposition where $\Lambda=\operatorname*{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$ with each $\lambda_{i}\geq 0$ and $V$ is the orthonormal matrix of eigenvectors of $B_{uc}^{T}B_{uc}$. Further, let $q=V^{T}\overline{u}_{uc}$. Then, $\|q\|_{\infty}\leq\|V^{T}\|_{\infty}\|\overline{u}_{uc}\|_{\infty}\leq\|V\|_{1}$, where we use the sub-multiplicative property of norms, $\|V^{T}\|_{\infty}=\|V\|_{1}$ and $\|\overline{u}_{uc}\|_{\infty}\leq 1$ for $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$. Changing the optimization variable to $q$ using this property,
@@
where the optimal $q^{*}$ has each component $q_{i}^{*}=\pm\|V\|_{1}$, maximizing the objective since each $\lambda_{i}>0$. Next, consider $T_{2}$. For any vector $\eta\in\mathbb{R}^{p}$, $\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\eta^{T}\overline{u}_{uc}=\|\eta\|_{1},$ where the optimal $\overline{u}_{uc}$ is given by $\overline{u}_{uc}^{*}=\operatorname*{sign}(\eta)$. Thus,
@@
Substituting (23), (24) and (25) in (22), the result follows.
Remark 1
@@
The bound obtained in (21) may be conservative. However, in Section 4, we consider the special case when $p=1$, i.e., when control authority is lost over a single actuator, and derive a closed-form expression for $\overline{E}_{M}(x_{0},t_{f})$. We then use that expression to derive bounds on the energetic resilience metric (9).
Remark 2
@@
We note some similarities between the results of this section and the work of Bouvier et al. (2021) on time-optimality for linear driftless systems. We have shown that the energy-optimal control signals in the nominal (14) and malfunctioning (19) cases are constant. Similarly, it was shown by Bouvier et al. (2021) that the time-optimal control signals in both the nominal and malfunctioning cases are constant; however, closed-form expressions for this signal are not provided by Bouvier et al. (2021), unlike in (14) and (19). We also note in the malfunctioning case that each component of the worst-case uncontrolled input $u_{uc}^{*}(t)$ takes on a constant value of either $+1$ or $-1$ in the interval $[0,t_{f}]$. A similar result in (Bouvier et al., 2021) shows that the worst-case uncontrolled input maximizing reachable time also has constant components with maximum allowable amplitude. However, closed-form expressions for this worst-case uncontrolled input are also not provided by Bouvier et al. (2021).
4 Resilience to the Loss of One Actuator
@@
We now consider the special case of losing control authority over one actuator, i.e., $p=1$. We derive a closed-form expression for the worst-case total energy (8) and the corresponding worst-case uncontrolled input. Using this expression, we derive a bound on the resilience metric (9).
4.1 Worst-case Total Energy
When $p=1$, $B_{uc}\in\mathbb{R}^{n}$ and $u_{uc}(t)\in\mathbb{R}$. From (8) and (20),
@@ 4.1 Worst-case Total Energy
$=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\Biggl{\{}t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}$
$+2x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}+\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t\Biggr{\}}.$ (26)
Recall that $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ implies $\|\overline{u}_{uc}\|_{\infty}\leq 1$, or $-1\leq\overline{u}_{uc}\leq 1$. Note that the second term in the supremum is maximized by $\overline{u}_{uc}^{*}=\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})$ as in Proposition 1, and reduces to $2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|$. When $\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})\in\{-1,+1\}$, this choice of $\overline{u}_{uc}^{*}$ also maximizes the other terms inside the supremum so that
@@
Alternatively, if $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}=0$, the second term in the supremum vanishes, reducing the value of $\overline{E}_{M}(x_{0},t_{f})$. The worst-case total energy is thus achieved only when $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\neq 0$, with the constant uncontrolled input
$u_{uc}^{*}(t)=\operatorname*{sign}(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0})~{}\text{ for all }~{}t\in[0,t_{f}].$ (28)
4.2 Energetic Resilience Metric
We now consider the energetic resilience metric $r_{M}(t_{f},R)$ in (9), and use the expression (27) to derive bounds on the this metric in the case when control authority is lost over a single actuator. We note that a bound on $r_{M}(t_{f},R)$ can be derived when control authority is lost over $p>1$ actuators using (21). However, this bound is more conservative since (21) is only an upper bound on the worst-case total energy and not an exact expression as in (27).
Proposition 2
When $p=1$, the energetic resilience metric $r_{M}(t_{f},R)$ is bounded from below as follows: